OpenStudy (anonymous):

What is the sum of the product and quotient of 9 and 3?

OpenStudy (anonymous):

The sum (+) of the Product(*) and the quotient (/) of 9 and 3 ? (a*b) = product (a/b) = quotient To sum them means to add them (a*b)+(a/b) I think this is what the question is asking, and you just have to replace values for a and b here with 9 and 3

OpenStudy (anonymous):

Here are the answer choices a. 27 b. 12 c. 42 d. 30

OpenStudy (anonymous):

well plugging in the values.. a =9 b =3 you get product + quotient (9*3)+(9/3) = ???